# Learning control variables and instruments for causal analysis in observational data

### v1

@@ 5 Simulation study
The intermediate rows present the results for a violation of Assumption 4, selection on observables, when setting $\delta=2$, $\gamma=0$. As the sample size increases, the detection rates of $Z$ (det.Z) and the confounder variables (det.X) both go to zero, as expected. Furthermore, the estimated violation $\hat{\theta}$ conditional on using $Z$ as an instrument (est) is now bounded away from zero across all samples. The lower rows of Table 1 provide the performance measures under a violation of Assumption 5, instrument validity, when considering $\delta=0$, $\gamma=0.5$. Again and as it should be the case, both the detection rates of $Z$ (det.Z) and the confounder variables (det.X) tend to zero as the sample size increases. At the same time, the estimated violation conditional on using $Z$ as an instrument (est) is substantially biased across all samples.
@@
Our simulations point to the consistency of our procedure in terms of selecting the correct instrument when Assumptions 4 and 5 are satisfied, and the rejection of the null hypothesis when violations of these assumptions occur. However and as a word of caution for empirical applications, our simulations also show that the sample size might need to be quite substantial for our testing approach to perform adequately, as the finite sample performance in our simulations is not convincing in moderate samples of just several thousand observations.
@@ 6 Empirical application
This section provides an empirical application to the Job Corps (JC) training program, one of the largest and most comprehensive job training programs for disadvantaged youth in the US that offers various types of academic and vocational training. The data come from the National Job Corps Study, a randomized experiment conducted in the mid-to-late 1990s in the United States to evaluate the effectiveness of JC on various labour market outcomes. Schochet et al. (2001) and Schochet et al. (2008) evaluate the impact of random program assignment on a wide range of labor market outcomes and find positive effects on education, employment, and earnings in the longer run.
We employ our testing methodology to analyze the experimental JC data as available in the causalweight package for the statistical software R, provided by Bodory and Huber (2018). This dataset named JC comprises 9240 observations and 46 variables. Among these variables is a binary indicator denoting participation in any form of academic or vocational training within the first year following program assignment, which serves as our treatment variable of interest $D$. It is noteworthy that while assignment to JC is randomized, actual participation in training might be selective due to non-compliance with the assignment. For instance, some individuals eligible for JC may choose not to participate. Our outcome variable are weekly earnings in the fourth year post-JC assignment. Alongside random program assignment, which stands as an obvious candidate instrument, our vector $Q$ encompasses 28 pre-assignment characteristics. These include gender, age, ethnicity, education level, mother tongue, marital status, a binary indicator for having at least one child, previous labor market participation and earnings, household size, parents’ education, welfare receipt during childhood, as well as variables related to health and health behavior. Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively. In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$. As in the simulations, we apply lasso regression for nuisance parameter estimation.
@@
When running our algorithm, it selects random program assignment as the instrument, yielding a very high p-value of 76% when running the conditional independence test. On the one hand, this indicates the validity of the instrumental variable (IV), implying not only the randomness of program assignment inherent to the experimental design. It also suggests that the assignment does not directly affect the earnings outcome other than through the treatment (e.g. through motivation or disappointment when being or not being eligible for the program). On the other hand, the testing result suggests that the selection-on-observables assumption holds for the treatment when controlling for the pre-assignment covariates available in the data. Consequently, given the covariates, our result suggests that we may evaluate the average treatment effect (ATE) on the total population. In contrast, employing an IV-based approach utilizing random assignment as the instrument for effect estimation would, under specific additional assumptions like treatment monotonicity in the instrument, only permit assessing the local average treatment effect (LATE) on the subpopulation of compliers, whose participation in training aligns with the random assignment, as discussed in Imbens and Angrist (1994) and Angrist et al. (1996).
@@ 7 Conclusion
In this paper, we introduced a machine learning-based algorithm based on a novel doubly robust score function designed to detect and test, in a data-adaptive manner, the presence of control variables sufficient for identifying treatment effects in observational data, as well as variables satisfying instrument validity. The method learns the partition of observed variables into instruments and control variables from the data, relying on the condition that valid instruments are conditionally independent of the outcome given the treatment and control variables that imply the conditional treatment exogeneity. This represents an important advancement over previously suggested tests of identifying assumptions for causal inference in observational data, which rely on a priori assumptions about whether a variable is a control variable or an instrument.
We demonstrated the consistency of the method for detecting control variables and instruments (if they exist) under certain regularity conditions and investigated its finite sample performance through a simulation study, which supports the consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations. Finally, we applied our algorithm to empirical labor market data from the Job Corps study. Our approach (as expected) selected random assignment into the Job Corps program as valid instrument, while also suggesting that training participation in the first year after assignment, our treatment of interest, is exogenous conditional on pre-assignment covariates.
@@ A.1 Proof of moment condition and Neyman orthogonality of $\tilde{\psi}$
$\quad\tilde{\psi}(W,\theta,\eta)$ (A.1)
@@
$\quad\tilde{\psi}_{1}(W,\theta,\eta)$
@@
$\quad\partial_{r}E[\tilde{\psi}_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
@@
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(Y-\mu_{0}(D,X,1))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\quad+2E\bigg{[}((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))$
$\quad\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\bigg{]}$
@@
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
@@
$\quad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,1)-\mu_{0}(D,X,1))\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
$\quad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,0)-\mu_{0}(D,X,0))\mathbb{E}\left[\frac{(1-Z)}{1-p_{0}(D,X)}\Bigg{|}D,X\right]\right]$
$\quad=0$
@@
$\quad\partial_{r}\mathbb{E}[\tilde{\psi}_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
@@
$\quad-\mathbb{E}\left[\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\quad+\mathbb{E}\left[\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\quad-\mathbb{E}\left[\frac{Z(Y-\mu_{0}(1,D,X))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\quad-\mathbb{E}\left[\frac{(1-Z)(Y-\mu_{0}(0,D,X))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]=0$

### v2

@@ 5 Simulation study
The intermediate rows of Table 1 and middle column of Figure 2 present the results for a violation of Assumption 4, selection on observables, when setting $\delta=2$, $\gamma=0$. As the sample size increases, the selection rates of $Z$ (sel.Z) and the confounder variables (sel.conf) both go towards zero, as expected. The rate of finding a nonempty set $\hat{\mathcal{P}}_{pass}$, which is equivalent the sum of the selection rates sel.Z, sel.noconf and sel.conf, also goes towards zero, finding no identification as expected. Furthermore, the estimated violation conditional on using $Z$ as an instrument ($\hat{\theta}$) is statistically different from zero under the larger sample sizes $n=4000,16000$. We note that in settings where Assumption 4 is violated, the selection rate of a confounder (sel.conf) in the binary scenario is still slightly high, around 20% in the largest sample for the binary instrument, athough this rate is 0% for the continuous instrument. [^11] In further experiments (not reported here), strengthening the degree of confounding by increasing the value of $\gamma$ leads to a substantial reduction in the rate at which a confounder is selected as the best candidate instrument. The lower rows of Table 1 and final column of Figure 2 provide the performance measures under a violation of Assumption 5, instrument validity, when considering $\delta=0$, $\gamma=0.5$. Again and as it should be the case, both the selection rates of $Z$ (sel.Z) and the confounder variables (sel.conf) tend to zero as the sample size increases. At the same time, the estimated violation conditional on using $Z$ as an instrument ($\hat{\theta}$) is large in absolute terms and highly statistically significant across all sample sizes, and $\hat{\mathcal{P}}_{pass}$ is empty 89% of the time for binary and 99% of the time for continuous instruments for the largest sample size.
We now consider a modification of the previous data generating process that entails the existence of multiple instruments. Specifically, the true set of instruments $(Z_{1}...Z_{3})$ now comprises the last three elements in $Q$, i.e. variables 8 to 10, which are constructed equivalently to $Z$ in the previous setup. We investigate the performance when Assumption 5 is violated for all of the instruments. This implies that the set of confounders remains the same as before ($X_{1}...X_{4}$), while the set of noise variables now only consists of three elements ($X_{5}...X_{7}$). Results for the multiple instrument simulations are shown in Table 2, and are similar to those shown in the single instrument results. In the case of multiple binary instruments and no violations of the assumptions, the selection rate of $Z$ (sel.Z) is up to 96% in the largest sample size. As in the single instrument scenario, under a violation of Assumption 4, selection on observables, the selection rates of the true instruments (sel.Z) and confounding variables (sel.conf) go to zero. For multiple continuous instruments (Panel B), results are consistent with the single instrument setting, with the exception of the slightly lower increase of the instrument selection rate in the larger sample size compared to the binary version (67% in the case of multiple instruments, compared to 64% under the single instrument scenario). However, in all settings of the multiple continuous instrument case, the rate of selecting a confounder (sel.conf) goes to zero and the estimate magnitudes and $\hat{\mathcal{P}}_{pass}$ behave as expected, moving towards zero and the empty set under no violation of the assumptions and away from zero otherwise.
Table 2: Simulations: Multiple Instruments
PANEL A: Binary Instruments
N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
1000 80% 1% 18% 0.002 0.008 0.007
4000 84% 2% 12% 0.000 0.002 0.002
16000 96% 0% 0% 0.000 0.000 0.000
Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
1000 86% 1% 11% 0.010 0.012 0.014
4000 40% 2% 40% 0.010 0.005 0.007
16000 0% 3% 16% 0.011 0.003 0.004
Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$ $\forall Z$)
1000 70% 0% 21% 0.053 0.035 0.050
4000 15% 2% 48% 0.058 0.018 0.026
16000 0% 1% 8% 0.058 0.009 0.013
PANEL B: Continuous Instruments
N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
1000 62% 1% 6% -0.010 0.035 0.031
4000 64% 1% 4% -0.003 0.017 0.016
16000 67% 2% 0% -0.001 0.010 0.008
Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
1000 38% 2% 2% -0.027 0.036 0.017
4000 14% 0% 1% -0.025 0.021 0.007
16000 0% 0% 0% -0.040 0.007 0.003
Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$ $\forall Z$)
1000 20% 2% 10% -0.116 0.090 0.037
4000 0% 2% 6% -0.142 0.064 0.019
16000 0% 2% 0% -0.140 0.017 0.010
Notes: columns $\hat{\theta}$’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z_{3}$ (the last element in $Q$) as instrument, its standard deviation, and the average of the standard errors across all samples, respectively. ‘sel.Z’ gives the share of simulations in which one of the instruments $Z_{1}$ to $Z_{3}$ is selected as best candidate instrument and simultaneously yields a p-value >0.30>0.30 (or 30%) when testing conditional independence based on the estimate of $\hat{\theta}$. ‘sel.noconf’ is an equivalent measure for covariates $X_{5}$ to $X_{7}$, corresponding to the share of selecting a non-confounder as best candidate instrument and having a p-value >> 30% when testing conditional independence. ‘sel.conf’ is an equivalent measure for covariates $X_{1}$ to $X_{4}$, corresponding to the share of selecting an observable confounder the best candidate instrument and having a p-value >> 30% when testing conditional independence.
@@
Our simulations point to the consistency of our procedure in terms of selecting the correct instrument when Assumptions 4 and 5 are satisfied, and the rejection of the null hypothesis when violations of these assumptions occur. Where there is more than one binary candidate instrument and both assumptions are satisfied, the procedure chooses one of the correct instruments $90\%$ of the time (sel.Z) with the largest sample size (N = 16000). However and as a word of caution for empirical applications, our simulations also show that the sample size might need to be quite substantial for our testing approach to perform adequately.
@@ 6 Empirical application
This section provides an empirical application to the Oregon Health Insurance Experiment (OHI), in which a group of low-income adults were selected randomly by lottery to be eligible to apply for Medicaid, a type of public health insurance in the United States. The experiment began in early 2008, when the state of Oregon drew names from the 90,000 individuals who had signed up for a limited number of spaces in the Medicaid program, and a rolling notification of participants regarding their lottery outcomes ran until October 2009. Previous work uses the random program assignment as an instrument for health insurance coverage and finds increases in primary, emergency, and preventive care utilisation, decreases in out-of-pocket medical expenditures, and other improvements in self-reported physical and mental health (Finkelstein et al., 2012; Baicker et al., 2013; Taubman et al., 2014; Finkelstein et al., 2019).
We apply our testing methodology to the experimental OHI data, which is publicly available from the NBER at https://www.nber.org/research/data/oregon-health-insurance-experiment-data, using the sample definition adopted by Finkelstein et al. (2012). The data comprises 23,762 observations and 692 variables. Among the variables is a binary indicator denoting whether an individual was ever enrolled in Medicaid following rolling notification of lottery selection by October 2009. This serves as the treatment variable of interest $D$. It is noteworthy that while assignment to OHI is randomized, actual enrolment in Medicaid might be selective due to non-compliance with the assignment. For instance, some individuals selected as eligible for Medicaid may choose not to enrol. We examine two outcome variables: the number of primary care visits and the number of prescription medications taken. Alongside random program assignment, which stands as an obvious candidate instrument, our vector $Q$ encompasses 218 pre-assignment characteristics. These include gender, age, ethnicity, education level, household size, employment and labour income, as well as a suite of variables capturing pre-treatment health status, healthcare utilisation and health expenditure. We also include individual indicator variables for whether there are any missing values for each pre-treatment characteristic.
Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with five folds ($K=5$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively. In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$. As in the simulations, we apply lasso regression for nuisance parameter estimation, although we also report estimates using random forest and gradient boosting as alternatives. Furthermore, we drop observations with extremely low or high propensity scores, which cause estimation instability and lead to high variance. For a discussion of such trimming based on the propensity score, see, for instance, Crump et al. (2009) and Lechner and Strittmatter (2019). Specifically, we discard observations for which the estimated propensity score $p_{l}(D,X)=P(Z\in Z_{l}\mid D,X)$, defined for a partition $Z_{l}$ of the candidate instrument (whether continuous or discrete), falls outside the interval $0.01<p_{l}(D,X)<0.99$. We require that no more than 5% of observations be trimmed using this rule in order for the candidate instrument to be considered.
Table 3: Empirical Application
PANEL A: Primary Care Visits
Variable $\hat{\theta}$ se p-value
Lasso 0.000 0.000 0.997
Random Forest -0.000 0.013 0.999
XGBoost -0.070 0.207 0.735
PANEL B: Number of Prescriptions
Method $\hat{\theta}$ se p-value
Lasso 0.001 0.000 0.148
Random Forest -0.007 0.010 0.472
XGBoost 0.073 0.088 0.407
Notes: this table reports the estimate $\hat{\theta}$ from equation (3.6) with five folds ($K=5$) when using the doubly robust score functions in equations (3.4) and (3.5) in the case of binary and continuous candidate instruments, respectively. ‘se‘ and ‘p-value‘ report the standard error and reported p-value for estimate $\hat{\theta}$ on the random program assignment variable.
Results are depicted in Table 3, where we show the estimate $\hat{\theta}$ for the random program assignment variable. Although there are multiple candidate instruments in $\hat{S}$, none of the other candidates were retained as possible IVs, as their instrument propensity score is either too large or too small to satisfy the trimming rule. [^12] We report every variable selected in the first stage to be included in $\hat{S}$, sorted by decreasing p-value, in Appendix Table A1, as well as the number of observations dropped following the trimming rule.
@@
When running our algorithm, it selects random program assignment as an instrument, yielding estimates of $\hat{\theta}$ with values very close to zero. According to our chosen p-value threshold of 30%, the random assignment is a valid instrument across all three algorithms for the doctor visits outcome. In the case of the lasso algorithm, the estimate is exactly zero and a p-value of 99.7% is reported, when running the conditional independence test. On the one hand, this indicates the validity of the instrumental variable (IV), implying not only the randomness of program assignment inherent to the experimental design. It also suggests that the assignment does not directly affect the earnings outcome other than through the treatment (e.g., through motivation or disappointment when being or not being eligible for the program). On the other hand, the testing result suggests that the selection-on-observables assumption holds for the treatment when controlling for the pre-assignment covariates available in the data. For the number of prescriptions outcome, the test results are less conclusive: the random forest algorithm yields a p-value of 0.472 and the XGBoost algorithm yields a p-value of 0.407, passing the 0.3 threshold, but the lasso results in a lower p-value. Consequently, given the covariates, our result suggests that we may evaluate the average treatment effect (ATE) on the total population for the doctor visits outcome. In contrast, employing an IV-based approach utilizing random assignment as the instrument for effect estimation would, under specific additional assumptions like treatment monotonicity in the instrument, only permit assessing the local average treatment effect (LATE) on the subpopulation of compliers, whose participation in training aligns with the random assignment, as discussed in Imbens and Angrist (1994) and Angrist et al. (1996).
@@ 7 Conclusion
In this paper, we introduced a machine learning-based algorithm based on a novel doubly robust score function designed to detect and test, in a data-adaptive manner, the presence of control variables sufficient for identifying treatment effects in observational data, as well as variables satisfying instrument validity. The algorithm may conclude that identification is given, enabling the estimation of the ATE as well as LATE or conditional average treatment effects (CATEs). The method learns the partition of observed variables into instruments and control variables from the data, relying on the condition that valid instruments are conditionally independent of the outcome given the treatment and control variables that imply the conditional treatment exogeneity. This represents an important advancement over previously suggested tests of identifying assumptions for causal inference in observational data, which rely on a priori assumptions about whether a variable is a control variable or an instrument.
We demonstrated the consistency of the method for detecting control variables and instruments (if they exist) under certain regularity conditions and investigated its finite sample performance through a simulation study, which supports the consistency result. However, as a word of caution for empirical applications, the results also suggest that decent performance of the test might require a relatively large dataset rather with more than just several 1000 observations. Further experiments (not reported in this work) improved the performance of the test in terms of reducing the rate at which a confounder is chosen as the candidate instrument, when the strength of violations of Assumptions 4 and 5 are increased by increasing the values of $\gamma$ and $\delta$. However, as the strength of these violations may not be so severe in practice, we maintain lower values for $\gamma$ and $\delta$ in our reported simulation results. Finally, we applied our algorithm to empirical health data from the Oregon Health Insurance Experiment. Our approach (as expected) always selected random assignment into the program as valid instrument, across a variety of algorithms for the nuisance function learner.
@@ A.1 Proof of moment condition and Neyman orthogonality of $\tilde{\psi}$
$\qquad\tilde{\psi}(W,\theta,\eta)$ (A.1)
@@
$\qquad\tilde{\psi}_{1}(W,\theta,\eta)$
@@
$\qquad\partial_{r}E[\tilde{\psi}_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
@@
$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(Y-\mu_{0}(D,X,1))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(Y-\mu_{0}(D,X,0))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]$
$\qquad+2E\bigg[((\mu(D,X,1)-\mu_{0}(D,X,1))-(\mu(D,X,0)-\mu_{0}(D,X,0)))$
$\qquad\left(\frac{(Y-\mu_{0}(D,X,1))\cdot Z}{p_{0}(D,X)}-\frac{(Y-\mu_{0}(D,X,0))\cdot(1-Z)}{1-p_{0}(D,X)}\right)\bigg]$
@@
$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
@@
$\qquad-2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,1)-\mu_{0}(D,X,1))\mathbb{E}\left[\frac{Z}{p_{0}(D,X)}\Bigg|D,X\right]\right]$
$\qquad+2E\left[(\mu_{0}(D,X,1)-\mu_{0}(D,X,0))(\mu(D,X,0)-\mu_{0}(D,X,0))\mathbb{E}\left[\frac{(1-Z)}{1-p_{0}(D,X)}\Bigg|D,X\right]\right]$
$\qquad=0$
@@
$\qquad\partial_{r}\mathbb{E}[\tilde{\psi}_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
@@
$\qquad-\mathbb{E}\left[\frac{Z(\mu(D,X,1)-\mu_{0}(D,X,1))}{p_{0}(D,X)}\right]$
$\qquad+\mathbb{E}\left[\frac{(1-Z)(\mu(D,X,0)-\mu_{0}(D,X,0))}{1-p_{0}(D,X)}\right]$
$\qquad-\mathbb{E}\left[\frac{Z(Y-\mu_{0}(1,D,X))(p(D,X)-p_{0}(D,X))}{p_{0}(D,X)^{2}}\right]$
$\qquad-\mathbb{E}\left[\frac{(1-Z)(Y-\mu_{0}(0,D,X))(p(D,X)-p_{0}(D,X))}{(1-p_{0}(D,X))^{2}}\right]=0$